# RCPU: Rotation-Constrained Error Compensation for Structured Pruning of a Large Language Model

### v1

@@ Abstract
In this paper, we propose a rotation-constrained compensation method to address the errors introduced by structured pruning of large language models (LLMs). LLMs are trained on massive datasets and accumulate rich semantic knowledge in their representation space. In contrast, pruning is typically carried out with only a small amount of calibration data, which makes output mismatches unavoidable. Although direct least-squares fitting can reduce such errors, it tends to overfit to the limited calibration set, destructively modifying pretrained weights. To overcome this difficulty, we update the pruned parameters under a rotation constraint. This constrained update preserves the geometry of output representations (i.e., norms and inner products) and simultaneously re-aligns the pruned subspace with the original outputs. Furthermore, in rotation-constrained compensation, removing components that strongly contribute to the principal directions of the output makes error recovery difficult. Since input dimensions with large variance strongly affect these principal directions, we design a variance-aware importance score that ensures such dimensions are preferentially kept in the pruned model. By combining this scoring rule with rotation-constrained updates, the proposed method effectively compensates errors while retaining the components likely to be more important in a geometry-preserving manner. In the experiments, we apply the proposed method to LLaMA-7B and evaluate it on WikiText-2 and multiple language understanding benchmarks. The results demonstrate consistently better perplexity and task accuracy compared with existing baselines.
@@ 1 Introduction
Figure 1: Overview of the proposed rotation-constrained compensation framework. Input activations are scored by a variance-aware importance score, less important columns are pruned, and the retained subspace is updated through the proposed rotation-constrained fitting. The pruned output vectors (arrows) are rotated to align with the original output vectors, showing how the proposed update compensates for pruning errors.
@@
Large language models (LLMs) are driving a rapid wave of transformation and are now being deployed across a wide range of applications, including code assistance, conversational agents, search and summarization, agentic execution, and content generation Jiang et al. (2025); Fan et al. (2024); Zhang et al. (2025); Wang et al. (2024). At the same time, their inference costs in computation and memory remain substantial, creating significant bottlenecks for deployment. In mobile and embedded settings, there is a strong need for model compression techniques that reduce computational cost while preserving task performance Saha & Xu (2025); Girija et al. (2025). Against this backdrop, a variety of efficient methods have been explored, such as quantization, knowledge distillation, and pruning Miao et al. (2025). Among them, structured pruning typically removes parameters at the granularity of weight-matrix rows or columns, and in some cases even entire transformer blocks Kim et al. (2024). Such structured removal directly reduces the parameter count and thereby lowers memory usage and inference cost He & Xiao (2024).
@@
Early work targeting LLMs, such as LLM-Pruner, demonstrated the feasibility of compression but tends to rely on downstream fine-tuning to recover high accuracy Ma et al. (2023). Therefore, methods that preserve accuracy without re-training are desirable. WANDA is widely used for unstructured sparsification, and prunes using simple activation-aware heuristics with only a small calibration set without downstream fine-tuning Sun et al. (2024). Its structured variant, Wanda-sp, provides a simple column-pruning importance score and serves as a familiar baseline in structured pruning An et al. (2024). Yet, removing columns inevitably introduces output discrepancies. In real deployments, the data available for calibration is often limited. Thus, pruning decisions must be made under sparse observations, and how errors are handled becomes decisive for overall performance. Recent work, FLAP, has shown that compensating the mean component of post-pruning errors with a bias term can be practical and effective An et al. (2024). Nonetheless, input-dependent directional mismatches are not easily addressed by a constant bias. As another possible approach, one might consider least-squares style fitting that directly minimizes output error. However, such a broad parameter update, under limited calibration data, risks overfitting that damages knowledge acquired during pretraining.
@@
Motivated by these limitations, in this paper, we propose RCPU, a Rotation-Constrained Parameter Update, to reduce pruning error while preserving the norm and inner-product structure of output representations. Figure 1 shows the overview of RCPU. Compared to general linear least-squares updates, restricting the update to rotations preserves angles and lengths, which helps avoid geometric distortion under small calibration sets. We formulate the alignment between the retained outputs and the original outputs as an Orthogonal Procrustes problem and, for each layer, estimate the optimal rotation and use it to update the parameters. The constraint reduces the update’s degrees of freedom, which improves statistical stability and makes it less prone to overfitting. Moreover, since the choice of retained components strongly affects the effectiveness of rotation-constrained compensation, we adopt a simple pruning score that augments weight magnitude and input scale with input variance. By doing this, components contributing to principal output directions are preferentially kept. We combine this scoring rule with rotation-constrained updates, and RCPU effectively compensates errors while retaining the components likely to be more important in a geometry-preserving manner. In the experiments, we apply RCPU to an existing LLM and evaluate it on a variety of language understanding benchmarks. As a result, we demonstrate consistent improvements over existing baselines in both perplexity and task accuracy.
The main contributions of this paper are as follows:
@@ 2.1 Notation and setup
We consider a linear sub-layer in a transformer block (e.g., attention output projection or MLP down-projection) with weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$. From a small calibration set of $N$ token positions, we record input activations $\mathbf{X}\in\mathbb{R}^{d_{\text{in}}\times N}$ and the corresponding original outputs $\mathbf{Y}=\mathbf{W}\mathbf{X}\in\mathbb{R}^{d_{\text{out}}\times N}$.
@@
To compensate the discrepancy without using dropped columns, we transform the kept output by a linear map $\mathbf{A}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{out}}}$ and seek $\mathbf{A}\mathbf{Z}\approx\mathbf{Y}$. Concretely, we consider minimizing
$\mathcal{L}(\mathbf{A})\;=\;\bigl\lVert\mathbf{Y}-\mathbf{A}\mathbf{Z}\bigr\rVert_{F}^{2}.$ (3)
2.2 Unconstrained least-squares compensation
A straightforward way to minimize the error defined in equation 3 is to apply a least-squares fitting, and $\mathbf{A}$ can be calculated by
$\mathbf{A}^{\star}\;=\;\arg\min_{\mathbf{A}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{out}}}}\;\lVert\mathbf{Y}-\mathbf{A}\mathbf{Z}\rVert_{F}^{2}.$ (4)
The normal equations give $\mathbf{A}^{\star}=\mathbf{Y}\mathbf{Z}^{\top}(\mathbf{Z}\mathbf{Z}^{\top})^{\dagger}$, where $(\cdot)^{\dagger}$ denotes the Moore–Penrose pseudoinverse.
@@ 2.2 Unconstrained least-squares compensation
(i) Geometric distortion: an unconstrained linear fit may introduce scaling and shear that reduce in-sample error while altering angles and norms in the output space, which can harm generalization beyond the calibration set. Desideratum: preserve the geometry of the outputs as much as possible. (ii) Overparameterization: a full $d_{\text{out}}\times d_{\text{out}}$ map has $d_{\text{out}}^{2}$ degrees of freedom; when the number of calibration tokens is small, the normal equations might become ill-conditioned and the estimate exhibits high variance. Desideratum: reduce the effective degrees of freedom to improve statistical stability. (iii) Undesired coupling: a original linear map mixes output coordinates, potentially degrading well-preserved directions to correct poorly preserved ones. Desideratum: prefer updates that act locally with respect to the kept subspace and maintain structure across output directions.
These issues motivate constraining the compensation map $\mathbf{A}$ to geometry-preserving transformations with limited degrees of freedom, while keeping the update within $\mathbf{W}_{K}$.
@@ 3 RCPU (Rotation Constrained Parameter Update)
We target the error introduced by structured column pruning in linear sub-layers of a transformer block. After pruning, the kept subspace still carries most of the signal, yet its orientation relative to the original outputs can be misaligned. Our idea is to re-align the orientation by an rotation-constrained parameter update computed from a small calibration set. Restricting the update to rotations preserves norm and inner-product relationships of output representations, helping reduce error while maintaining the pretrained geometry.
@@ 3.1 Rotation-based compensation via Orthogonal Procrustes
Equation 5 is known as the classical Orthogonal Procrustes problem Golub & Van Loan (2013). It corresponds to the least-squares formulation in equation 4, but with the compensation map $\mathbf{A}$ constrained to an rotation matrix $\mathbf{Q}$.
@@
In other words, applying the update makes the new output $\widetilde{\mathbf{W}}_{K}\mathbf{X}_{K}=\mathbf{Q}^{\star}\mathbf{Z}$. Thus, the kept component $\mathbf{Z}$ is explicitly rotated by $\mathbf{Q}^{\star}$ so that its orientation matches the original output $\mathbf{Y}$. Finally, we replace the sub-layer weight with the compact matrix $\widehat{\mathbf{W}}:=\widetilde{\mathbf{W}}_{K}\in\mathbb{R}^{d_{\text{out}}\times k}$, meaning that columns in $D$ are physically removed.
@@
$(\mathbf{Q}^{\star},s^{\star})\;=\;\arg\min_{\mathbf{Q}^{\top}\mathbf{Q}=\mathbf{I},\,s\geq 0}\;\bigl\lVert\mathbf{Y}-s\,\mathbf{Q}\mathbf{Z}\bigr\rVert_{F}^{2}.$ (8)
@@
Pruning removes the $\mathbf{W}_{D}\mathbf{X}_{D}$ term in equation 1, but the kept subspace often still captures much of the useful signal. By restricting the update to rotation (with an optional isotropic scaling), the retained subspace can be re-aligned with the original output geometry while preserving angles and relative norms. This avoids the arbitrary scaling and shear distortions that an unconstrained least-squares fit may introduce under limited calibration.
@@ 3.3 Algorithm and complexity
We apply the procedure layerwise to a designated set of linear sub-layers. Algorithm 1 summarizes the steps. This procedure is greedy and layerwise: each layer–sub-layer Procrustes subproblem admits a closed-form global minimizer ( equation 6 and equation 9), but the overall routine is not a joint global optimization across the network.
@@
  • $\mathbf{X}_{K}=\mathbf{X}_{[K,:]}$
  • $\mathbf{W}_{K}=\mathbf{W}_{[:,K]}$
  • $\mathbf{Z}=\mathbf{W}_{K}\mathbf{X}_{K}$
@@
Complexity. Per treated sub-layer, computing scores takes $O(d_{\text{in}}(d_{\text{out}}+N))$, forming $\mathbf{Z}$ costs $O(d_{\text{out}}kN)$, and constructing $\mathbf{M}=\mathbf{Y}\mathbf{Z}^{\top}$ requires $O(d_{\text{out}}^{2}N)$. The dominant cost is the SVD of $\mathbf{M}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{out}}}$, which is typically $O(d_{\text{out}}^{3})$. Thus the overall complexity is cubic in $d_{\text{out}}$, on par with unstructured pruning methods such as SparseGPT Frantar & Alistarh (2023).
@@ 4.1 Settings
We evaluate RCPU on LLaMA-7B as the base model Touvron et al. (2023).
As baselines, we use WANDA-sp, a pruning method based on weight magnitude and activation scale; and FLAP, a bias-based error compensation method.
Following prior work, we use WikiText-2 as the calibration dataset with 128 samples Merity et al. (2016). We perform pruning based on the input channels of o_proj and down_proj, and simultaneously remove the corresponding positions in the other projection matrices. For the attention modules, we prune at the head level. Parameter updates, however, are applied only to o_proj and down_proj. We evaluate pruning ratios of 10%, 20%, and 30%.
@@ 4.2 Perplexity
Table 1: Representative PPL of WANDA-sp, FLAP, and proposed methods under pruning. The best score in each setting is highlighted in bold, while the second-best score is underlined. See also Figure 2.
Method Pruning Ratio PPL ($\downarrow$)
Llama-7B (Original) 0% 12.4
WANDA-sp 10% 14.66
FLAP 10% 14.14
RCPU (Rot.) 10% 13.55
RCPU (Rot.+Scale) 10% 13.52
WANDA-sp 20% 16.70
FLAP 20% 15.36
RCPU (Rot.) 20% 14.40
RCPU (Rot.+Scale) 20% 14.55
WANDA-sp 30% 24.13
FLAP 30% 18.59
RCPU (Rot.) 30% 18.35
RCPU (Rot.+Scale) 30% 18.21
(a) WANDA-sp score-based pruning with under different update methods.
(b) Proposed score-based pruning under different update methods.
Figure 2: Perplexity versus pruning ratio for different pruning scores and update methods.
Table 1 summarizes perplexity on WikiText-2 across pruning ratios. As we can see from Table 1, RCPU consistently improves upon WANDA-sp and performs on par with or better than FLAP. These results show that rotation-constrained updates can be competitive with bias-based correction, despite offering higher degrees of freedom. While the advantage is not uniform at all pruning levels, the geometry-preserving nature of rotational transformations helps prevent the distortions that often arise from unconstrained updates.
@@
We also compare the scaled variant (Rot.+Scale) to plain rotation. The results show only marginal differences, indicating that introducing a global rescaling does not significantly alter perplexity. We believe this is because the rotation already aligns the retained subspace with the dominant output directions, effectively discarding less informative components. As a result, the overall output norm is well preserved, and an additional scaling factor brings little further benefit.
Figure 2 focuses on the effect of different error-compensation methods (LS or Rot.). Compared to no compensation, rotation-constrained updates consistently suppress the growth of perplexity, and the benefit becomes more pronounced at higher pruning ratios. In contrast, least-squares fitting provides little to no improvement, and in many cases even worsens perplexity, reflecting its tendency to overfit under limited calibration. Another notable trend is that within both WANDA-sp-based and variance-aware scores, rotation compensation is consistently the most effective. Considering that the only difference between the two scores is the inclusion of an input-variance factor, the results suggest that variance-aware scoring further strengthens orientation recovery by retaining columns that align with dominant output directions. Taken together, these findings indicate that beyond the choice of compensation method, the proposed variance-aware score itself offers a stronger foundation for structured pruning, yielding lower perplexity across pruning ratios compared to WANDA-sp-based scoring.

### v2

# RCPU: Rotation-Constrained Error Compensation for Structured Pruning of Large Language Models

@@ Abstract
In this paper, we propose a rotation-constrained compensation method to address the errors introduced by structured pruning of large language models (LLMs). LLMs are trained on massive datasets and accumulate rich semantic knowledge in their representation space. In contrast, pruning is typically carried out with only a small amount of calibration data, which makes output mismatches unavoidable. Although direct least-squares fitting can reduce such errors, it tends to overfit to the limited calibration set, destructively modifying pretrained weights. To overcome this difficulty, we update the pruned parameters under a rotation constraint. This constrained update preserves the geometry of output representations (i.e., norms and inner products) and simultaneously re-aligns the pruned subspace with the original outputs. Furthermore, in rotation-constrained compensation, removing components that strongly contribute to the principal directions of the output makes error recovery difficult. Since input dimensions with large variance strongly affect these principal directions, we design a variance-aware importance score that ensures such dimensions are preferentially kept in the pruned model. By combining this scoring rule with rotation-constrained updates, the proposed method effectively compensates errors while retaining the components likely to be more important in a geometry-preserving manner. In the experiments, we apply the proposed method to Llama-7B and Llama-2-13B, and evaluate it on WikiText2 and multiple language understanding benchmarks. The results demonstrate consistently better perplexity and task accuracy compared with existing baselines.
@@ 1 Introduction
Figure 1: Overview of RCPU framework. Input activations are scored by a variance-aware importance score, less important columns are pruned, and the retained subspace is updated through the rotation-constrained fitting. The pruned output vectors (arrows) are rotated to align with the original output vectors, showing how RCPU compensates for pruning errors.
@@
Large language models (LLMs) are driving a rapid wave of transformation and are now being deployed across a wide range of applications, including code assistance, conversational agents, search and summarization, agentic execution, and content generation Jiang et al. (2025); Fan et al. (2024); Zhang et al. (2025); Wang et al. (2024). At the same time, their inference costs in computation and memory remain substantial, creating significant bottlenecks for deployment. In mobile and embedded settings, there is a strong need for model compression techniques that reduce computational cost while preserving task performance Saha and Xu (2025); Girija et al. (2025). Against this backdrop, a variety of efficient methods have been explored, such as quantization, knowledge distillation, and pruning Miao et al. (2025). Among them, structured pruning typically removes parameters at the granularity of weight-matrix rows or columns, and in some cases even entire transformer blocks Kim et al. (2024). Such structured removal directly reduces the parameter count and thereby lowers memory usage and inference cost He and Xiao (2024).
@@
Early work targeting LLMs, such as LLM-Pruner, demonstrated the feasibility of compression but tends to rely on downstream fine-tuning to recover high accuracy Ma et al. (2023). Therefore, methods that preserve accuracy without re-training are desirable. WANDA is widely used for unstructured sparsification, and prunes using simple activation-aware heuristics with only a small calibration set without downstream fine-tuning Sun et al. (2024). Its structured variant, Wanda-sp, provides a simple column-pruning importance score and serves as a familiar baseline in structured pruning An et al. (2024). Yet, removing columns inevitably introduces output discrepancies. In real deployments, the data available for calibration is often limited. Thus, pruning decisions must be made under sparse observations, and how errors are handled becomes decisive for overall performance. Recent work, FLAP, has shown that compensating the mean component of post-pruning errors with a bias term can be practical and effective An et al. (2024). Nonetheless, input-dependent directional mismatches are not easily addressed by a constant bias. As another possible approach, one might consider least-squares style fitting that directly minimizes output error. However, such a broad parameter update, under limited calibration data, risks overfitting that damages knowledge acquired during pretraining even if using regularization method like Ridge Hoerl and Kennard (1970).
@@
Motivated by these limitations, in this paper, we propose RCPU, a Rotation-Constrained Parameter Update, to reduce pruning error while preserving the norm and inner-product structure of output representations. Figure 1 shows the overview of RCPU. Compared to general linear least-squares updates, restricting the update to rotations preserves angles and lengths, which helps avoid geometric distortion under small calibration sets. We formulate the alignment between the retained outputs and the original outputs as an Orthogonal Procrustes problem and, for each layer, estimate the optimal rotation and use it to update the parameters. The constraint reduces the update’s degrees of freedom, which improves statistical stability and makes it less prone to overfitting. Moreover, since the choice of retained components strongly affects the effectiveness of rotation-constrained compensation, we adopt a simple pruning score that augments weight magnitude and input scale with input variance. By doing this, components contributing to principal output directions are preferentially kept. We combine this scoring rule with rotation-constrained updates, and RCPU effectively compensates errors while retaining the components likely to be more important in a geometry-preserving manner. In the experiments, we apply RCPU to existing LLMs and evaluate it on a variety of language understanding benchmarks. As a result, we demonstrate improvements over existing baselines in both perplexity and task accuracy. The main contributions of this paper are as follows:
@@ 2.1 Notation and setup
We consider a linear sub-layer in a transformer block (e.g., attention output projection or multi layer perceptron (MLP) down-projection) with weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$. From a small calibration set of $N$ token positions, we record input activations $\mathbf{X}\in\mathbb{R}^{d_{\text{in}}\times N}$ and the corresponding original outputs $\mathbf{Y}=\mathbf{W}\mathbf{X}\in\mathbb{R}^{d_{\text{out}}\times N}$.
@@
To compensate the discrepancy without using dropped columns, one possible approach is to update the kept parameters. Concretely, the following regularized optimization problem can be considered:
$\mathcal{L}(\mathbf{W^{\star}})=\lVert\mathbf{Y}-\mathbf{W}^{\star}\mathbf{X}_{K}\rVert_{F}^{2}+\lambda\lVert\mathbf{W}^{\star}-\mathbf{W}_{K}\rVert_{F}^{2},$ (3)
where $\lambda$ is regularization hyper-parameter. This corresponds to Ridge regression, in which the updated weights are penalized for deviating from the original ones.
2.2 Least-squares compensation
A straightforward way to minimize the error defined in equation 3 is to apply a least-squares fitting. The closed-form solution is
$\mathbf{W}_{K}^{\star}=(\mathbf{Y}\mathbf{X}_{K}^{\top}+\lambda\mathbf{W}_{K})\,(\mathbf{X}_{K}\mathbf{X}_{K}^{\top}+\lambda\mathbf{I})^{-1}.$ (4)
@@ 2.2 Least-squares compensation
(i) Geometric distortion: an unconstrained linear fit may introduce scaling and shear that reduce in-sample error while altering angles and norms in the output space, which can harm generalization beyond the calibration set. Desideratum: preserve the geometry of the outputs as much as possible. (ii) Limited effectiveness of regularization: The $\lambda$ values that minimize calibration perplexity do not stabilize the estimator and, in our experiments, often degrade downstream performance. Desideratum: ensure stability in a way that does not depend on regularization tuned purely for calibration set.
These issues motivate restricting the compensation update to geometry-preserving transformations with limited flexibility, while ensuring that the update operates only within the kept subspace.
@@ 3 RCPU (Rotation Constrained Parameter Update)
We target the error introduced by structured column pruning in linear sub-layers of a transformer block. After pruning, the kept subspace still carries most of the signal, yet its orientation relative to the original outputs can be misaligned. Our idea is to re-align the orientation by a rotation-constrained parameter update computed from a small calibration set. Restricting the update to rotations preserves norm and inner-product relationships of output representations, helping reduce error while maintaining the pretrained geometry.
@@ 3.1 Rotation-based compensation via Orthogonal Procrustes
Equation 5 is known as the classical Orthogonal Procrustes problem Golub and Van Loan (2013). It corresponds to the least-squares formulation in equation 3 and equation 4, but with the compensation update restricted to an orthogonal matrix.
@@
In other words, applying the update makes the new output $\widetilde{\mathbf{W}}_{K}\mathbf{X}_{K}=\mathbf{Q}^{\star}\mathbf{Z}$. Thus, the kept component $\mathbf{Z}$ is explicitly rotated by $\mathbf{Q}^{\star}$ so that its orientation matches the original output $\mathbf{Y}$. Finally, we replace the sub-layer weight with the compact matrix $\widehat{\mathbf{W}}:=\widetilde{\mathbf{W}}_{K}\in\mathbb{R}^{d_{\text{out}}\times k}$, meaning that columns in $D$ are physically removed. Note that we follow common practice and simply refer to $Q$ as a “rotation” throughout the paper, even though it may occasionally include a reflection.
@@
$(\mathbf{Q}^{\star},s^{\star})\;=\;\arg\min_{\mathbf{Q}^{\top}\mathbf{Q}=\mathbf{I},\,s>0}\;\bigl\lVert\mathbf{Y}-s\,\mathbf{Q}\mathbf{Z}\bigr\rVert_{F}^{2}.$ (8)
@@
Pruning removes the $\mathbf{W}_{D}\mathbf{X}_{D}$ term in equation 1, but the kept subspace often still captures much of the useful signal. By restricting the update to rotation (with an optional isotropic scaling), the retained subspace can be re-aligned with the original output geometry while preserving angles and relative norms. This avoids the arbitrary scaling and shear distortions that the least-squares fit may introduce under limited calibration.
@@ 3.3 Algorithm and complexity
We apply the procedure layerwise to a designated set of linear sub-layers. Algorithm 1 summarizes the steps. This procedure is greedy and layerwise: each layer Procrustes subproblem admits a closed-form minimizer ( equation 6 and equation 9), but the overall routine is not a joint global optimization across the network.
@@
  • $\mathbf{X}_{K}=\mathbf{X}_{[K,:]}$, $\mathbf{W}_{K}=\mathbf{W}_{[:,K]}$, and $\mathbf{Z}=\mathbf{W}_{K}\mathbf{X}_{K}$
@@
Complexity. Per treated sub-layer, computing scores takes $O(d_{\text{in}}(d_{\text{out}}+N))$, forming $\mathbf{Z}$ costs $O(d_{\text{out}}kN)$, and constructing $\mathbf{M}=\mathbf{Y}\mathbf{Z}^{\top}$ requires $O(d_{\text{out}}^{2}N)$. The dominant cost is the SVD of $\mathbf{M}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{out}}}$, which is typically $O(d_{\text{out}}^{3})$. Thus the overall complexity is cubic in $d_{\text{out}}$, on par with unstructured pruning methods such as SparseGPT Frantar and Alistarh (2023).
@@ 4.1 Settings
We evaluate RCPU on Llama-7B and Llama-2-13B as the base models Touvron et al. (2023a; b)  [^1] Codes are available at https://github.com/harutaro/rcpu.. As baselines, we use WANDA-sp, a pruning method based on weight magnitude and activation scale; and FLAP, a bias-based error compensation method. In addition, we compare RCPU with SliceGPT. Since SliceGPT does not support Llama-1-7B, we evaluate it on Llama-2-13B with benchmarks.
Following prior work, we use WikiText-2 as the calibration dataset Merity et al. (2016). We perform pruning based on the input channels of o_proj and down_proj, and simultaneously remove the corresponding positions in the other projection matrices. For the attention modules, we prune at the head level. Parameter updates, however, are applied only to o_proj and down_proj. We evaluate pruning ratios of 10%, 20%, and 30%.
@@ 4.2 Perplexity
Figure 2: Perplexity versus calibration-set size for RCPU on Llama-7B.
@@
Table 1: Representative PPL ($\downarrow$) of WANDA-sp, FLAP, and RCPU under $N_{\mathrm{calib}}=128,512$. The best score in each setting is in bold, while the second-best score is underlined. See also Figure 3.
Method PR Llama-7B Llama-2-13B
128 512 128 512
Original 0% 12.4 12.4 10.98 10.98
WANDA-sp 10% 14.66 14.53 12.29 12.29
FLAP 10% 14.14 14.11 12.22 12.08
RCPU (Rot.) 10% 13.55 13.48 11.65 11.57
RCPU (Rot.+Scale) 10% 13.52 13.45 11.66 11.57
WANDA-sp 20% 16.70 16.96 14.62 14.62
FLAP 20% 15.36 15.07 14.49 14.14
RCPU (Rot.) 20% 14.40 14.83 13.12 12.75
RCPU(Rot.+Scale) 20% 14.55 14.81 13.07 12.72
WANDA-sp 30% 24.13 26.20 61.66 63.35
FLAP 30% 18.59 18.31 17.15 16.71
RCPU (Rot.) 30% 18.35 16.99 16.99 16.01
RCPU (Rot.+Scale) 30% 18.21 16.91 16.88 15.96
(a) WANDA-sp score+ comp. ($N_{\mathrm{calib}}=$128)
(b) Proposed score + compensation ($N_{\mathrm{calib}}=$128)
(c) WANDA-sp score+ comp. ($N_{\mathrm{calib}}=$512)
(d) Proposed score + compensation ($N_{\mathrm{calib}}=$512)
Figure 3: PPL vs P.R. ratio for different calibration sizes and compensation methods on Llama-7B.
First of all, we conducted experiments regarding PPL, which changes $N_{\mathrm{calib}}$ (The number of calibration samples). Figure 2 reports how the PPL of RCPU varies with the number of calibration samples. We observe that PPL drops rapidly as $N_{\mathrm{calib}}$ increases and becomes roughly stable once $N_{\mathrm{calib}}$ reaches around 64. Based on this trend, we adopt $N_{\mathrm{calib}}=$ 128 and 512 as the calibration set sizes for other experiments. While 128 is a common choice in prior work, we additionally include 512, which lies well within the empirically stable region, providing more reliable evaluations.
@@
Table 1 summarizes PPL on WikiText-2 across pruning ratios. As we can see from Table 1, RCPU consistently improves upon WANDA-sp and performs better than FLAP regardless of the number of calibration samples and models. These results show that rotation-constrained updates can be competitive with bias-based compensation. While the advantage is not uniform at all pruning levels, the geometry-preserving nature of rotational transformations helps prevent the distortions that often arise from unconstrained updates. We also compare the scaled variant (Rot.+Scale) to plain rotation. The results show only marginal differences, indicating that introducing a global rescaling does not significantly alter PPL. We believe this is because the rotation already aligns the retained subspace with the dominant output directions, effectively discarding less informative components. As a result, the overall output norm is well preserved, and an additional scaling factor brings limited benefit.
Next, in Figure 3, we compare the compensation effectiveness of RCPU and LS (Least Square)+Ridge regularization (equation 4). We performed a grid search over $\lambda\in\{10^{-6},10^{-5},\ldots,10^{6}\}$ and reported the result on best $\lambda$ (See Section 7.1). Figure 3 focuses on the effect of different error-compensation methods (LS or Rot.) in different calibration sizes. From Figure 3(a) and Figure 3(b), we observe that with $N_{\mathrm{calib}}=128$, the proposed rotation-based compensation achieves the best PPL. In Figure 3(a), the regularized LS even worsens PPL, reflecting its tendency to overfit under limited calibration. In contrast, Figure 3(c) and Figure 3(d) show that when $N_{\mathrm{calib}}=512$, both rotation and Ridge-based compensation effectively reduce PPL, but the regularized LS update contributes more strongly to the improvement. Similar trend is observed in Llama-2-13B as shown in Appendix 7.2. Although it is intuitively expected that the effect of the least square fitting becomes larger as the number of calibration samples increases, we emphasize that this does not necessarily translate into better downstream benchmark performance. Indeed, for example in Table 6 in Appendix, LS-based compensations are not the top performer, whereas RCPU often achieves the best accuracy. According to Hastie et al. (2001), the degree of freedom in ridge regression is computed by $d_{\mathrm{out}}\sum_{i}\frac{\sigma_{i}^{2}}{\sigma_{i}^{2}+\lambda}$, where $\sigma$ denotes the singular value of the input matrix. Using this equation and the best $\lambda$, we obtain the degree of freedom values in the range $1.395\times 10^{9}$ to $1.578\times 10^{9}$ across each pruning ratio. In contrast, the degree of freedom in RCPU is given by $\frac{d_{\mathrm{out}}(d_{\mathrm{out}}-1)}{2}$ since $Q$ is constrained to the orthogonal matrix. Its degree of freedom is $5.36\times 10^{8}$, which is smaller than that of LS+Ridge. This implies that, from the standpoint of preserving the pretrained knowledge of LLMs, the rotation-based compensation tends to be more robust. We also highlight that, in the context of LLMs, selecting an appropriate regularization hyper-parameter $\lambda$ can be computationally expensive, as it requires repeatedly computing large matrix inverses for multiple candidate values of $\lambda$. In contrast, our method has no hyper-parameters, avoiding this overhead.